1. The price of a unit of Good 1 is Tk20. The price of a unit of Good 2   is Tk5. The consumer's income is Tk100. Graph the consumer's budget constraint (with Good 2 on the horizontal axis.) Indicate the intercepts and slope.
1
Expert's answer
2020-10-02T06:54:28-0400

The consumer's budget constraint (with Good 2 on the horizontal axis) is a line, which intersects vertical axis at 100/20 = 5 units of Good 1 and horizontal axis at 100/5 = 20 units of Good 2.

The budget constraint equation is 20Q1 + 5Q2 = 100, or Q1 = 5 - 0.25Q2, so the slope is -0.25.